A plank of length 2.89 m is to be balanced on a see-saw. On one side of the plank hangs a mass 1 =6.32 kg, and on the other side of the plank there is a massless, frictionless pulley. This pulley has a mass 2​​ =22.2 kg on one side and a mass 3 =6.25 kg on the other side. What is the distance from the see-saw's pivot to the point where mass 1 must hang in order for this plank to be stationary after being released from rest?
1
Expert's answer
2020-03-06T10:24:53-0500

As per the given data in the question,

Length of the plank =2.89m

The hanged mass on one side of the plank (m1) =6.32kg

Mass hanged on one side of the pulley (m2) =22.2 Kg

Mass hanged on other side of the pulley(m3) =6.25 kg

Let the acceleration of the blocks, which hanged over the pulley =a

The hanged distance =?

Now, force on the pulley,

"m_2 g-T=m_2 a"

"\Rightarrow 22.2 g-T=22.2 a--------(i)"

Now,

"T-m_3g=m_3 a"

"T-6.25 g=6.25 a--------(ii)"

Adding equation (i) and (ii)


"\Rightarrow 22.2g-6.25g=22.2a+6.25a"


"\Rightarrow 15.95g=28.45a"


"\Rightarrow a=\dfrac{15.95}{28.45}"


"\Rightarrow a=0.56 m/sec^2"

So, tension on the string "(T)=6.25a+6.25g =64.75N" =

Hence total force on the plank "=2T=64.75\times2=129.5N"

Now, let the block is hanged from the point of rotation,

"m_1g\times (2.89-x)=129.5\times x"

"\Rightarrow 129.5\times x+m_1 g\times x=m_1g\times 2.89"

"\Rightarrow x(129.5+6.32)=6.32\times 9.8\times 2.89"

"\Rightarrow x(135.82)=178.995"

"\Rightarrow x=\dfrac{178.995}{135.82}m"

"x=1.32m"

from the side of pulley.